OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1).

OpenStudy (anonymous):

y - y1 = m(x - x1) m is slope y1 and x1 stand for x and y values on coordinate(x, y) Find slope. \[m = \frac{ y2 - y1 }{ x2-x1 }\]\[m = \frac{ 1 - 2 }{ 2 - (-3) } = \frac{ 1 - 2 }{ 2 + 3 }=\frac{ -1 }{ 5 }\]\[m = \frac{ -1 }{ 5 }\]So you have a slope of -1/5 and a given point (-3, 2). Plug the values into the equation. y - 2 = -1/5(x - (-3)) Two negatives(--) equal a positive(+) y - 2 = -1/5(x + 3)